### v1

# MAME: Multidimensional Adaptive Metamer Exploration with Human Perceptual Feedback

Abstract
Alignment between human brain networks and artificial models has become an active research area in both machine learning and neuroscience. A widely adopted approach to explore the functional alignment between them is to identify “metamers” for both humans and models. Metamers refer to input stimuli that are physically different but equivalent within a given system. If a model’s metameric space completely matched the human metameric space, the model would achieve functional alignment with humans. However, conventional methods lack a direct approach to searching for the human metameric space. As a result, researchers have had to first develop biologically inspired models and then infer about human metamers indirectly by testing whether model metamers also appear as metamers to humans. Here, we propose the Multidimensional Adaptive Metamer Exploration (MAME) framework, which enables direct high-dimensional exploration of human metameric space. MAME leverages online image generation guided by human perceptual feedback, allowing for a more flexible search process that mitigates the exploration constraint of human metamers. Specifically, the MAME framework modulates reference images across multiple dimensions by leveraging hierarchical responses from convolutional neural networks (CNNs). The generated metameric images are presented to human participants, and their perceptual discriminability from the reference images is assessed through a behavioral task. Based on participants’ responses, subsequent image generation parameters are adaptively updated online. Using our MAME framework, we successfully measured a human metameric space of over fifty dimensions within a single experiment. Experimental results showed that human discrimination sensitivity was lower for metameric images based on low-level features compared to high-level features, which image contrast metrics could not explain. The finding suggests that the model computes low-level information that is not essential for human perception. Our MAME framework has the potential to contribute to developing interpretable AI and to advance our understanding of brain function in neuroscience.
Keywords: cognitive neuroscience; vision science; model metamers; texture synthesis; psychophysics
1 Introduction
1.1 Functional Alignment Between Human Vision and Artificial Neural Networks
Recent advancements in artificial neural networks (ANNs) have not only improved their ability to solve visual tasks but have also positioned them as valuable analogies for understanding human visual processes. In particular, deep convolutional neural networks (CNNs) have been noted for their structural and functional similarities to the hierarchical processing observed in biological vision. These models have demonstrated considerable success in approximating the information flow from the primary visual cortex (V1) to higher visual areas such as V4 and the inferotemporal (IT) cortex in the ventral visual stream (Yamins and DiCarlo,, 2016). However, numerous studies have reported significant divergences between the behavioral responses of humans and ANN models (Geirhos et al., 2018a, ; Geirhos et al., 2018b, ; Szegedy,, 2013). Consequently, extensive research efforts have been directed toward achieving a functional alignment between human vision and artificial models (Geirhos et al.,, 2021).
1.2 Metamers as a tool for functional understanding of cognitive neural process
One widely adopted approach for functionally aligning models with humans is the identification of "metamers." Metamers refer to sets of input stimuli that are physically different yet processed as equivalent by a given system, such as humans or models. Here, consider an unknown system whose functional properties we aim to uncover. If the system produces the same output for different sets of physical inputs, i.e., metamers, it implies the existence of common factors to which the system is sensitive. Identifying these common factors would provide insight into what the system functionally computes.
Through this metamer-based approach, various visual functions have been uncovered in the history of vision sciences. A classic example is the finding that human color vision is mediated by three types of sensors (reviewed by Gegenfurtner, (2003)). The discovery that humans do not directly sense light as a function of wavelength was based on experimental results showing that physically different spectral compositions of light can appear identical to human observers, i.e., they form metamers.
Figure 1: The MAME framework assumes that images can be generated at arbitrary locations starting from a reference stimulus.
Furthermore, this approach has also been used to investigate hierarchical information processing in the brain. In the context of texture information processing, researchers have explored the features critical for texture perception by examining metamers and linking them to hierarchical processing in the brain (Julesz,, 1962; Bergen and Adelson,, 1988; Freeman and Simoncelli,, 2013; Freeman et al.,, 2013; Okazawa et al.,, 2015; Ziemba et al.,, 2024). For example, Freeman et al. (Freeman and Simoncelli,, 2013; Freeman et al.,, 2013) identified a model capable of generating metamers at mid-level stages of visual processing, such as V2, using texture synthesis algorithms. The exploration of metamers has been applied not only to texture perception but also to other visual processes, such as peripheral vision (Rosenholtz,, 2016; Wallis et al.,, 2019) and visual search (Rosenholtz et al.,, 2012), as well as to other sensory modalities, including audition (McDermott and Simoncelli,, 2011) and haptics (Kuroki et al.,, 2021).
The approach of using metamers for functional understanding has further advanced with the use of deep neural network models as artificial systems. Gatys et al., (2015) proposed an image generation technique that utilizes intermediate features in convolutional neural networks, summarizing them with statistical representations such as Gram matrices. This method has been particularly beneficial for cognitive neuroscience researchers interested in higher-order hierarchical visual processing. For example, in the context of object recognition research, a metamer-based approach combining this image algorithm with fMRI measurements has demonstrated that object-selective regions in higher ventral visual areas exhibit texture-like representation, computed by Gram matrices (Jagadeesh and Gardner,, 2022).
When exploring human metamers using image generation algorithms from deep learning models, one first generates model metamers from noise images for a specific target image and then tests whether humans perceive the generated images as equivalent to the target (Figure 1, left). In this case, the model used for metamer generation should closely align with human vision. In fact, adversarially-trained CNNs (Salman et al.,, 2020), which are known as robust for various tasks like humans, can generate metamers that better explain human peripheral vision (Harrington and Deza,, 2021). Furthermore, there have been attempts to evaluate the alignment between humans and models by assessing the extent to which metamers generated by models align with human metamers (Feather et al.,, 2019, 2023).
1.3 Challenges in Exploring High-Dimensional Human Metameric Spaces
A key limitation of the conventional metamer approach (Figure 1, left) is that the exploration of human metamers is indirect. First, a biologically plausible model must be prepared, followed by the generation of model metamers. Then, it is necessary to test whether these model metamers are also perceived as metamers by humans. This method can completely reveal the structure of the human metameric space only if a model that already aligns with human perception is available beforehand.
Instead, could we adopt a more direct approach to measuring human metamers? If such an approach were possible, we could first measure the human metameric space and then design models that align with it, ultimately achieving human-model functional alignment. In this case, rather than generating images from noise using a model, as in conventional studies, we could directly measure human metamers by freely exploring the human metameric space from a reference stimulus, as illustrated in Figure 1 (right). Furthermore, directly exploring human metamers is itself crucial for understanding human information processing, as exemplified by the discovery of color vision sensors.
However, this free exploration also presents significant challenges. The human metameric space is highly multidimensional, making it difficult to determine a clear search direction. Additionally, in offline measurement settings, the computational cost of such an exploration becomes extremely high.
Figure 2: The MAME framework creates tasks adaptively online and efficiently explores metamers.
  • It can explore more than 50 dimensions of human metameric space simultaneously in one psychophysical experiment.
  • Using a guiding CNN, we found that identifying the boundary of human metameric space guided by low-level features of the model was more difficult than when guided by high-level features. This finding suggests a weaker alignment between low-level processing in humans and the model, highlighting a potential guideline for designing models that better align with human perception.
  • The experimental environment is highly flexible because online image generation is performed on a Python server using a GPU, while its human experiment is conducted via a web browser with JavaScript. Specifically, it can be easily applied to crowdsourcing or fMRI experiments that need online feedback.
1.4 Proposed Approach: The MAME Framework
In this study, we propose a framework for exploring high-dimensional human metameric spaces directly, called Multidimensional Adaptive Metamer Exploration (MAME) (Figure 2). The MAME framework enables direct multidimensional exploration of the human metameric space by utilizing online image generation based on human feedback, thereby relaxing the constraints that a human-aligned model is needed beforehand. Specifically, in the MAME framework, the model is used solely to determine the search direction for human metamers Figure 1 (right). While it is better for the model to be biologically plausible when determining the direction, the final human metameric boundary can be determined independent of the model’s metamers. This aspect relaxes the constraints on human metamer exploration. Furthermore, to enhance the efficiency of multidimensional exploration, we employ online image generation guided by trial-wise human feedback, enabling the adaptive search for metameric boundaries.
In the experiment, we tested our MAME framework using a biologically plausible CNN, an adversarially trained ResNet50, on a natural image dataset, ImageNet. The key contributions of our MAME framework can be summarized as follows:
2 Methods
2.1 The MAME Framework
The MAME framework is designed to adaptively explore the boundaries of the human metameric space. This exploration is achieved by systematically perturbing a given reference image in multiple dimensions Fig. (1, right). The execution of the MAME framework relies on the following requirements:
  1. 1.
    Definition of the exploration direction of human metameric space The human metameric space, initially defined in an input image space $\mathbb{R}^{W\times H\times C}$, must be compressed to a dimension practical for exploration (on the order of tens of dimensions) and defined using a coordinate system. Using CNNs or biologically plausible models coupled with compression methods, one can define the subspace direction to explore human metameric boundaries.
  2. 2.
    Online Stimulus Generation After determining the direction of exploration, an image-generation algorithm is required that is capable of generating stimuli at any position within the defined metameric space. The stimulus generation must be completed within a time frame that ensures compatibility with the human feedback process in the MAME framework. For instance, when incorporated into an ABX test (Freeman and Simoncelli,, 2011), this means generation within 2–3 seconds.
  3. 3.
    Feedback from Tasks performed by Humans or Animals Finally, our MAME framework requires feedback from tasks performed by humans or animals. As described in the Introduction, metamers refer to images that are physically different but equivalent within a given system. However, the definition of “equivalence” changes across studies. Some adopt categorical equivalence based on behavioral responses (Feather et al.,, 2019), while others define it in terms of perceptual appearance (Freeman and Simoncelli,, 2011; Deza et al.,, 2017) using ABX tasks. In general, measuring perceptual appearance psychophysically is more time-consuming than categorical judgment. Moreover, evaluating the equivalence of neural or BOLD activations in neurophysiological or fMRI studies requires substantial costs. Our method can be applied to feedback from any type of task; however, it is particularly effective for tasks where response acquisition takes longer, as online image generation allows for a more efficient increase in the number of trials.
In the present study, the MAME framework was implemented by (1) defining the direction of human metameric space using a CNN and ICA, (2) implementing a stimulus generation method that perturbs the reference stimulus via gradient descent, and (3) incorporating ABX tasks as feedback.
2.2 Definition of the exploration direction of human metameric space
The definition of the metameric space used in this study is illustrated in Fig. 3. This definition is based on a CNN and independent component analysis (ICA). First, natural images were passed through a pre-trained CNN, while extracting Gram matrices from the activations of each layer as position-invariant features. Here, we used Gram matrices because the present study explored metamers based on the equivalence of perceptual appearance and focused summary statistics like textures representations (Jagadeesh and Gardner,, 2022). For the model, we used a robust variant of ResNet50 (He et al.,, 2016) (ResNet50 robustness ImageNet L2-Norm $\epsilon=3.0$) (Engstrom et al.,, 2019). This model was trained to be resistant to adversarial attacks and is known for having internal representations that are biologically plausible (Feather et al.,, 2023; Gaziv et al.,, 2024).
Next, ICA was performed on these Gram matrices, and components with high explanatory power were selected, thereby considering images as points in the ICA coordinate space. The exploration direction was defined along the axes of ICA from the reference stimulus.
Figure 3: Metameric space was defined using a CNN and ICA.
In the experiment, the metameric space was explored in a 54-dimensional space, determined as follows. Feature extraction layers in the CNN included conv1, layer3, and avgpool to explore low-, middle-, and high-level features. For ICA, the top three components from each layer were selected. Each component was explored in both positive and negative directions. Additionally, three conditions for stimulus presentation eccentricities (4°, 8°, and 12°) were considered. Consequently, the total dimensionality was calculated as 54 dimensions (3 x 3 x 3 x 2).
2.2.1 Image Dataset
The ImageNet validation dataset (Deng et al.,, 2009) was used as a set of natural images. Each image was first resized to 256×256 pixels, and then the central 224×224 pixels were cropped for use to match the input image set and image size used during the training of the pre-trained model.
2.2.2 Feature Extraction
Gram matrices were extracted from each layer of the robust ResNet50 model as features of the input images. A Gram matrix represents the similarity of activations across multiple filters, summed over spatial positions. Given the activation of the $l$-th layer of a CNN for an input image as $F_{jk}$, where $j$ denotes the filter index and $k$ denotes the spatial position, the Gram matrix $G_{ij}$ is defined as:
$G_{ij}=\sum_{k}F_{ik}F_{jk}.$
In this study, Gram matrices were extracted from multiple layers of the model using 1,000 images randomly selected from the ImageNet validation dataset. Since Gram matrices discard spatial information, they are well-suited for extracting spatially invariant features of images (Gatys et al.,, 2015, 2016).
2.2.3 Independent Component Analysis (ICA)
ICA was employed in this study to map the gram matrix feature onto a coordinate space. ICA assumes non-Gaussianity of signals and computes a linear transformation $W$ via gradient descent to separate source signals $X$ into independent components $S$ as $S=WX$. Each independent component (hereafter referred to as an ICA component) is standardized to have zero mean and a covariance matrix equal to the identity matrix.
In this study, ICA was applied to the dataset as follows. The Gram matrices of the target layers were vectorized to form feature vectors $x$ for each image. Using the feature vector x of 1,000 images, an input data matrix $X$ was constructed, and ICA was applied to obtain the independent component matrix:
$S=WX,\quad\mathbf{s}=W\mathbf{x},\quad X=SA^{\mathsf{T}}$
where $A$ represents the mixing matrix. The number of independent components was pre-set to 100, and the top 3 components with the highest contribution were selected. The contribution of each component was quantified using the explained variance (EV) defined as:
$\mathrm{EV}_{i}=1-\frac{\|X-\hat{X}_{i}\|_{\text{F}}^{2}}{\|X\|_{\text{F}}^{2}},$
where $\hat{X}_{i}$ is defined as:
$\hat{X}_{i}=S_{i}A^{\mathsf{T}},$
with $S_{i}$ being the matrix retaining only the $i$-th independent component. Unlike PCA, the total explained variance for multiple ICA components does not necessarily match the sum of the individual component contributions. The contribution of each component is shown in Fig. 4.
Figure 4: Explained variance of ICA components sorted in descending order. (Top-left) conv1, (Top-right) layer1, (Middle-left) layer2, (Middle-right) layer3, (Bottom-left) layer4, (Bottom-right) avgpool.
2.2.4 Selection of the Reference Stimuli Used in the Experiment
In this study, we focused on metamers of images located near the zero point of the ICA components for low-, mid-, and high-level features within the natural image distribution. The reason for focusing on them is to sample many reference images for our behavioral experiment so that participants do not memorize individual images.
From 50,000 images in the ImageNet validation dataset, we sampled images for three target layers whose ICA component distances from zero were in the lowest 20%. Among these, 614 images that were selected across all three layers were used as experimental stimuli.
Table 1: ICA component criteria and the number of selected images
Criterion* Layer Threshold Number of Img
Lowest 20% conv1 1.2194 614
layer3 1.7333
avgpool 0.7942
* An image satisfies a given criterion if its Euclidean distance in each layer is below the corresponding threshold.
2.3 Generation of Target Images
This section describes the method for generating images at designated positions using gradient descent. The image modification process was performed by specifying the target layer (conv1, layer3, and avgpool), the ICA component of interest (components 1, 2, 3), the direction of change (positive or negative), and the magnitude of change (a positive scalar corresponding to the target value).
The image generation procedure is as follows:
  1. 1.
    Compute the ICA components of the reference image for the target layer as $\mathbf{y}_{\text{original}}=W\mathbf{x}_{\text{original}}$, where $\mathbf{x}_{\text{original}}$ is the vectorized gram matrix of the target layer.
  2. 2.
    Define the target component as $\mathrm{comp}$ and the target value as $t$. The modified ICA component $\mathbf{y}_{\text{target}}$ is given by:
    $\mathbf{y}_{\text{target}}=\mathbf{y}_{\text{original}}+t\mathbf{e}_{\text{comp}},$
    where $\mathbf{e}_{i}$ is a unit vector with 1 at the $i$-th component and 0 elsewhere. This operation modifies only the target component while keeping all others unchanged. If the change direction is negative, the unit vector is multiplied by $-1$.
  3. 3.
    In each iteration, compute the ICA components of the current image as $\mathbf{y}_{\text{current}}=W\mathbf{x}_{\text{current}}$.
  4. 4.
    Compute the mean squared error (MSE) between $\mathbf{y}_{\text{target}}$ and $\mathbf{y}_{\text{current}}$ as the loss function.
  5. 5.
    Perform gradient descent using the Adam optimizer to minimize the loss and update pixel values of the current image while keeping the CNN model parameters fixed.
Examples of image generation for the three layers and three components used in the experiment are shown in Fig.5. The eccentricity condition did not affect image generation; the same algorithm was used to generate images, which were then presented at different locations as a four-degree square image.
Figure 5: An example of the images used in the experiment, generated with three layers and three components each. The numbers (in red and blue) below the layer names represent the target ICA values used during image generation.
2.4 Feedback from Behavioral task
The experiment consisted of an ABX test comparing generated images with reference images. Based on the participants’ response data, each image was updated until the exploration of human metamer boundaries converged.
2.4.1 Participants and Experimental Environment
Eight participants took part in the experiment. All experiments were approved by the Ethics Committee of the University of Tokyo and conducted in accordance with the guidelines of the Declaration of Helsinki. Written informed consent was obtained from all participants.
The MAME framework was implemented using a Python web framework (Django) and a browser-based experimental environment with JavaScript (Adolphe et al.,, 2022). Image generation was performed on a server machine with an NVIDIA RTX A2000 GPU, and the generated images were presented in the Google Chrome browser via a local server.
The experiment was conducted in a dark room, with participants’ heads fixed at a distance of 70 cm from the monitor. In all tasks, participants were instructed to fixate on a fixation point (Thaler et al.,, 2013) at the center of the monitor (Fig. 6).
Furthermore, to ensure that participants were indeed fixating on the designated point during the experiment, gaze tracking was performed using the Tobii Pro Spark (Tobii, Sweden). Trials in which fixation was not maintained were not used for parameter updates.
2.4.2 ABX Task
The procedure of the ABX task is as follows (Fig. 6). In each trial, participants are presented with three stimuli in sequence: stimulus A, stimulus B, and stimulus X. The first two stimuli (A and B) are the reference and generated images in a random order. The third stimulus (X) is a test stimulus that is the same as A or B. Participants are required to identify whether X is the same as A or B.
The presentation position was 4°, 8°, or 12° from the center depending on the eccentricity condition. The size of the presentation image was 4° x 4° in the visual angle.
Figure 6: Overview of the ABX test. A 200 ms stimulus presentation was alternated with a 500 ms blank period.
2.5 Exploration Using the Staircase Method
Based on the results of each ABX test, the target value was updated using the staircase method (2-up 1-down rule). Specifically, if a participant correctly answered two consecutive trials in the ABX task, the boundary of the human metameric space was narrowed by one step. Conversely, if the participant made an error in a single trial, the boundary was shifted in the direction of expansion. The threshold was estimated by averaging five reversals between correct and incorrect responses for each condition.
3 Results
The MAME framework was applied to each subject, and their metameric boundaries were obtained as ICA component values. Furthermore, analyses based on the RMS contrast of images and human contrast sensitivity were conducted, allowing the metameric boundaries to be computed using these indices.
3.1 Metameric Boundaries in ICA Components
The metameric boundaries represented by ICA components, obtained directly from the experiment, are shown in Table 2, Figure 7 and Figure 9. In Table 2 and Figure 7, the conditions of ICA components and direction were averaged within each participant; then, the mean thresholds were averaged across different participants. The horizontal axis of Figure 7 shows the layer conditions, and three eccentricities are plotted in each layer.
Across all subjects, the threshold values exhibited a logarithmic scale dependency on the layer. Specifically, the mean thresholds of conv1 were the highest in conv1, followed by layer3, and lowest in avgpool, decreasing as the layer depth increased. These results indicate that the boundaries of human metameric space are broader in the direction suggested by the robust CNN for lower-level features. This finding suggests a discrepancy in functional alignment between the model and human perception.
In contrast, the differences in thresholds due to variations in ICA components and direction were relatively minor compared to the inter-layer differences  2. This suggests that the human metameric space has uniform boundaries across different dimensional directions of the ICA space.
The results for eccentricity showed a gradual trend across conditions, with a particularly noticeable tendency for thresholds to increase as eccentricity increased in conv1 and layer3. To evaluate this trend in detail in terms of sensitivity to image contrast, the next section analyzes the results using RMS contrast
Figure 7: ICA threshold values at the metameric boundary. The mean and standard deviation across all subjects ($n=8$) are shown. The Conditional Min and Conditional Max in the figure represent the threshold search range, which depends on the step size and initial value for each layer. The upper bound is defined as the target threshold upper limit minus 0.4 times the step size, while the lower bound is the target threshold lower limit plus 0.4 times the step size.
Table 2: Mean threshold values computed by averaging across ICA components, search directions, and subjects for each layer and eccentricity. The standard deviation represents inter-subject variability ($n=8$).
Layer Eccentricity (°) Mean Std
conv1 4 47.915 8.8166
8 63.0404 11.8613
12 63.7918 8.5024
layer3 4 0.486 0.1275
8 0.5983 0.0686
12 0.6407 0.0533
avgpool 4 0.1538 0.0182
8 0.153 0.0253
12 0.1688 0.0203
3.2 Metameric Boundaries Based on RMS Contrast
The metameric boundaries obtained from the experiment were represented using the RMS contrast values of grayscale difference images. The RMS contrast is an index that quantifies the overall contrast strength of a luminance image and is defined as:
$C_{\mathrm{RMS}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(I_{i}-\langle I\rangle)^{2}}$
where $N$ denotes the number of pixels, $I_{i}$ represents the luminance value of each pixel, and $\langle I\rangle$ is the mean luminance of the image.
In this analysis, for each layer and eccentricity, grayscale difference images were generated by subtracting the reference image from the perturbed image using the target ICA component values at the metameric boundary (Figure 8a). The RMS contrast of these difference images was then computed.
The RMS contrast values exhibited a linear relationship with the ICA target values (Figures 8b, c, d). When visualizing the RMS contrast at the metameric boundary, it was observed that the values were particularly high in conv1 (Figure 8e). This suggests that, when measured using RMS contrast, human perception appears to be less sensitive to changes in conv1 compared to other layers.
Furthermore, the trend in eccentricity became more pronounced when analyzed using RMS contrast. Specifically, the slope of the threshold as a function of eccentricity was steepest in conv1, followed by layer3 and then avgpool, where it was more gradual.
Figure 8: (a) Examples of modulated images for each layer (displayed in grayscale). The target value $t$ determined based on the threshold for each layer. The rightmost column shows the difference images. (b, c, d) Relationship between ICA target values and the resulting RMS contrast for each layer (b: conv1, c: layer3, d: avgpool). For each target value, calculations were performed with $n=30$, randomly selecting ICA components and search directions. Error bars indicate standard deviation among difference images. (e) Mean and standard deviation of thresholds measured using RMS contrast. The average RMS contrast corresponding to the mean threshold values in Table 2 is shown, with standard deviation ($n=30$).
4 Discussion
4.1 Summary of Main Findings
The present study proposed the MAME framework, an experimental framework to directly explore human metamers using online feedback by humans with weak supervision of artificial models. The effectiveness of this framework was demonstrated by successfully estimating 54-dimensional image metamer boundaries. Experimental results indicated that, among image generation layers, the threshold in conv1 was higher based on the magnitude of ICA thresholds and RMS contrast, suggesting that changes in this layer were less noticeable compared to others.
4.2 Alignments of Higher and Lower Visual Processing in Humans and Models
In our MAME framework, it is important to note that when an image is shifted even slightly in the model’s exploration direction from the reference image, the shifted image is no longer a model metamer for the model. This is because, in the MAME framework, exploration is conducted along specific directions in the model’s activation space. As a result, a single step in this space exceeds the model’s metameric boundary in the image dimension, as illustrated in Figure 1 (right). This notion suggests that in the present exploration, the lower the mean threshold for humans at a given hierarchical level, the closer the boundaries of human and model metamers, indicating a stronger alignment between human and model representations at that level. In this regard, the low threshold observed in avgpool suggests a closer functional alignment between the metameric spaces of humans and the model.
This finding is consistent with recent studies using robust CNNs (Feather et al.,, 2019, 2023; Gaziv et al.,, 2024). For instance, it has been shown that adversarial attacks on high-level activations can induce category modulations that are also perceptible to humans (Gaziv et al.,, 2024). Furthermore, the alignment of high-level features has been a topic of intense discussion in both visual and LLM literature. For example, the Platonic Representation Hypothesis (Huh et al.,, 2024) proposes that ’neural networks, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces.’ From this perspective, the strong alignment of high-level representations between humans and modern neural models, trained with various data, might be a straightforward consequence.
However, our experimental results also suggest a functional misalignment between models and humans in low-level processing. Given that convolutional neural networks (CNNs) are originally inspired by the information processing of simple and complex cells in the primary visual cortex (Fukushima,, 1980), this result may seem counterintuitive. Moreover, visualizing the first layer of a pre-trained CNN reveals the emergence of Gabor-like filters. However, a crucial factor is that feedforward models such as ResNet exhibit strong scale dependency in their information processing. The first-layer filters in ResNet50 have a high-resolution size of 7×7 pixels and lack low-frequency components. Additionally, while the correlation between filters is known to be important for visual processing (Freeman and Simoncelli,, 2011; Okazawa et al.,, 2015), not all correlations among early-layer features in CNNs (i.e., the Gram matrix) might be necessarily relevant. Some attempts to address such early-stage mismatches have been reported to contribute to the development of more generalizable models (Dapello et al.,, 2020). Therefore, based on our findings, reconsidering the early-stage architecture of models might be crucial for improving functional alignment between humans and models in the future.
4.3 Limitations and Broader Future Applications
The main limitation of our study is that the explored stimulus conditions in our experiment do not comprehensively cover all possible directions of human metamer space in terms of clarifying human functional processing. While We used Gram matrices to define metameric equivalence, many other possible conditions remain unexamined. For instance, metamers can be generated not only based on Gram matrices but also using activation differences. Conducting complementary experiments with such alternative conditions will be essential for further advancing our understanding of brain function.
Finally, it is noteworthy to discuss the generalizability of the proposed MAME framework. MAME is a versatile framework that extends beyond metamer exploration, as it allows for multidimensional modulation starting from a reference stimulus. The concept of providing feedback based on model-defined multidimensional features is analogous to Decoded Neurofeedback (Shibata et al.,, 2011), making it a promising candidate for various applications.
Moreover, since the framework can be executed in a web browser, it is free from hardware constraints. It can be applied even in environments with strong magnetic fields, such as fMRI, and is also suitable for crowdsourcing experiments. These advantages position MAME as a foundational tool for a wide range of applied research.