### v2

# MAME: Multidimensional Adaptive Metamer Exploration with Human Perceptual Feedback

Abstract
Alignment between human brain networks and artificial models has become an active research area in vision science and machine learning. A widely adopted approach is identifying "metamers," stimuli physically different yet perceptually equivalent within a system. However, conventional methods lack a direct approach to searching for the human metameric space. Instead, researchers first develop biologically inspired models and then infer about human metamers indirectly by testing whether model metamers also appear as metamers to humans. Here, we propose the Multidimensional Adaptive Metamer Exploration (MAME) framework, enabling direct, high-dimensional exploration of human metameric spaces through online image generation guided by human perceptual feedback. MAME modulates reference images across multiple dimensions based on hierarchical neural network responses, adaptively updating generation parameters according to participants’ perceptual discriminability. Using MAME, we successfully measured multidimensional human metameric spaces within a single psychophysical experiment. Experimental results using a biologically plausible CNN model showed that human discrimination sensitivity was lower for metameric images based on low-level features compared to high-level features, which image contrast metrics could not explain. The finding suggests a relatively worse alignment between the metameric spaces of humans and the CNN model for low-level processing compared to high-level processing. Counterintuitively, given recent discussions on alignment at higher representational levels, our results highlight the importance of early visual computations in shaping biologically plausible models. Our MAME framework can serve as a future scientific tool for directly investigating the functional organization of human vision.
Keywords: metamers; texture synthesis; psychophysics; cognitive neuroscience; machine learning
1 Introduction
1.1 Functional Alignment Between Human Vision and Artificial Neural Networks
Recent advancements in artificial neural networks (ANNs) have not only improved their ability to solve visual tasks but have also positioned them as valuable analogies for understanding human visual processes. In particular, deep convolutional neural networks (CNNs) have been noted for their structural and functional similarities to the hierarchical processing observed in biological vision. These models have demonstrated considerable success in approximating the information flow from the primary visual cortex to higher visual areas in the ventral visual stream (Yamins and DiCarlo,, 2016). However, numerous studies have reported significant divergences between the behavioral responses of humans and ANN models (Geirhos et al., 2018a, ; Geirhos et al., 2018b, ; Szegedy,, 2013). Consequently, extensive research efforts have been directed toward achieving a functional alignment between human vision and artificial models (Geirhos et al.,, 2021).
1.2 Metamer exploration as a tool for functional understanding of cognitive neural processes
One widely adopted approach for functionally aligning models with humans is the identification of "metamers." Metamers refer to sets of input stimuli that are physically different yet processed as equivalent by a given system, such as humans or models. Here, consider an unknown system whose functional properties we aim to uncover. If the system produces the same output for different sets of physical inputs, i.e., metamers, it implies the existence of common factors to which the system is sensitive. Identifying these common factors would provide insight into what the system functionally computes.
Through this metamer-based approach, various visual functions have been uncovered in the history of vision sciences. A classic example is the finding that human color vision is mediated by three types of sensors (reviewed by (Gegenfurtner,, 2003)). The discovery that humans do not directly sense light as a function of wavelength was based on experimental results showing that physically different spectral compositions of light can appear identical to human observers, i.e., they form metamers.
Furthermore, this approach has also been used to investigate hierarchical information processing in the brain. In the context of texture information processing, researchers have explored the features critical for texture perception by examining metamers and linking them to hierarchical processing in the brain (Julesz,, 1962; Bergen and Adelson,, 1988; Freeman and Simoncelli,, 2013; Freeman et al.,, 2013; Okazawa et al.,, 2015; Ziemba et al.,, 2024). For example, Freeman et al. (Freeman and Simoncelli,, 2013; Freeman et al.,, 2013) identified a model capable of generating metamers at mid-level stages of visual processing, such as V2, using texture synthesis algorithms. The exploration of metamers has been applied not only to texture perception but also to other visual processes, such as peripheral vision (Rosenholtz,, 2016; Wallis et al.,, 2019) and visual search (Rosenholtz et al.,, 2012), as well as to other sensory modalities, including audition (McDermott and Simoncelli,, 2011) and haptics (Kuroki et al.,, 2021).
The approach of using metamers for functional understanding has further advanced with the use of deep neural network models (DNNs) as artificial systems. Gatys et al. (Gatys et al.,, 2015) proposed an image generation technique that utilizes intermediate features in convolutional neural networks, summarizing them with statistical representations such as Gram matrices. This method has been particularly beneficial for cognitive neuroscience researchers interested in higher-order hierarchical visual processing. For example, in the context of object recognition research, a metamer-based approach combining this image algorithm with fMRI measurements has demonstrated that object-selective regions in higher ventral visual areas exhibit texture-like representation, computed by Gram matrices (Jagadeesh and Gardner,, 2022). Furthermore, there have been attempts to evaluate the alignment between humans and models by assessing the extent to which metamers generated by DNN models align with human metamers (Feather et al.,, 2019, 2023).
1.3 Challenges in Exploring High-Dimensional Human Metameric Spaces
A key limitation of approaches that use deep neural networks (DNNs) to explore human metamers is that these approaches are inherently indirect. This limitation becomes evident when examining how previous studies have explored human metamers (Figure 1).
Figure 1: A conceptual illustration of the MAME framework. In previous studies, human metamers were explored indirectly by generating model metamers from noise images and presenting them to human observers. In contrast, our approach proposes a method for directly measuring human metamers by performing multidimensional exploration starting from a reference image.
In previous studies, model metamers are first synthesized from noise to match a target image in a given DNN, and human observers are then asked whether these images appear equivalent to the original target (Figure 1, left). In this approach, the effectiveness of human metamer exploration critically depends on how well the model used for synthesis aligns with human visual perception. Indeed, adversarially trained convolutional neural networks (CNNs), which are known to exhibit human-like robustness across various tasks (Salman et al.,, 2020), have been shown to generate metamers that better account for human peripheral vision (Harrington and Deza,, 2021).
However, this method can fully reveal the structure of human metameric representations only if a suitably aligned model is available in advance. In this sense, human metamer exploration remains an indirect approach.
Instead, could we adopt a more direct approach to measuring human metamers? If such an approach were possible, we could first measure the human metameric space and then design models that align with it, ultimately achieving human-model functional alignment. In this case, rather than generating images from noise using a model, as in conventional studies, we could directly measure human metamers by freely exploring the human metameric space from a reference stimulus, as illustrated in Figure 1 (right). Furthermore, directly exploring human metamers is itself crucial for understanding human information processing, as exemplified by the discovery of color vision sensors.
However, this free exploration also presents significant challenges. The human metameric space is highly multidimensional, making it difficult to determine a clear search direction. Additionally, in offline measurement settings, the computational cost of such an exploration becomes extremely high.
1.4 Proposed Approach: The MAME Framework
In this study, we propose a framework for exploring high-dimensional human metameric spaces directly, called Multidimensional Adaptive Metamer Exploration (MAME) (Figure 2). The MAME framework enables direct multidimensional exploration of the human metameric space by utilizing online image generation based on human feedback, thereby relaxing the constraints that a human-aligned model is needed beforehand. Specifically, in the MAME framework, the model is used solely to determine the search direction for human metamers. While it is better for the model to be biologically plausible when determining the direction, the final human metameric boundary can be determined independent of the model’s metamers. This aspect relaxes the constraints on human metamer exploration. Furthermore, to enhance the efficiency of multidimensional exploration, we employ online image generation guided by trial-wise human feedback, enabling the adaptive search for metameric boundaries.
Figure 2: Exploring human metamers using the MAME framework.The exploration space is defined by the model (in this study, ICA components of CNN activations). Stimuli are synthesized along the defined directions, and their perceptual equivalence to a reference image is evaluated via a human behavioral task. Based on the feedback, the estimated human metamer is adaptively updated, and the task is repeated until convergence.
In the experiment, we tested our MAME framework using a biologically plausible CNN, an adversarially trained ResNet50, on a natural image dataset, ImageNet. The key contributions of our MAME framework can be summarized as follows:
  • It can explore multidimensional human metameric space simultaneously in one psychophysical experiment.
  • Using a guiding CNN, we found that identifying the boundary of human metameric space guided by low-level features of the model was more difficult than when guided by high-level features. This finding suggests a weaker alignment between low-level processing in humans and the model, highlighting a potential guideline for designing models that better align with human perception.
  • The experimental environment is highly flexible because online image generation is performed on a Python server using a GPU, while the human experiment is conducted via a web browser with JavaScript.
2 Methods
2.1 The MAME Framework
The MAME framework is designed to adaptively explore the boundaries of the human metameric space. This exploration is achieved by systematically perturbing a given reference image in multiple dimensions (Fig. 1, right). The execution of the MAME framework relies on the following requirements:
  1. 1.
    Definition of the exploration direction of human metameric space The human metameric space, initially defined in an input image space $\mathbb{R}^{W\times H\times C}$, must be compressed to a dimension practical for exploration (on the order of tens of dimensions) and defined using a coordinate system. Using CNNs or biologically plausible models coupled with compression methods, one can define the subspace direction to explore human metameric boundaries.
  2. 2.
    Online Stimulus Generation After determining the direction of exploration, an image-generation algorithm is required that is capable of generating stimuli at any position within the defined metameric space. The stimulus generation must be completed within a time frame that ensures compatibility with the human feedback process in the MAME framework. For instance, when incorporated into an ABX test (Freeman and Simoncelli,, 2011), this means generation within 2–3 seconds.
  3. 3.
    Feedback from Tasks performed by Humans or Animals Finally, our MAME framework requires feedback from tasks performed by humans or animals. As described in the Introduction, metamers refer to images that are physically different but equivalent within a given system. However, the definition of “equivalence” changes across studies. Some adopt categorical equivalence based on behavioral responses (Feather et al.,, 2019), while others define it in terms of perceptual appearance (Freeman and Simoncelli,, 2011; Deza et al.,, 2017) using ABX tasks. Our method can be applied to feedback from any type of task if the computation speed of image generation is fast enough for real-time online experiments.
Depending on the research objective, different methods can be selected for each requirement. In the present study, our aim was to demonstrate the effectiveness of MAME, a novel comprehensive framework, under specific conditions. To this end, we evaluated its effectiveness using a canonical approach within the context of metamer exploration research. Specifically, the MAME framework was implemented by (1) defining the direction of human metameric space using a CNN and ICA, (2) implementing a stimulus generation method that perturbs the reference stimulus via gradient descent, and (3) incorporating ABX tasks as feedback. The specific motivations for selecting each method are described in detail in the corresponding sections of the Methods.
2.2 Definition of the exploration direction of human metameric space
In this section, we describe how the metameric search space was defined in the present study. We constructed the search space based on the activations of a robust variant of ResNet50 (He et al.,, 2016) (ResNet50 robustness ImageNet L2-Norm $\epsilon=3.0$) (Engstrom et al.,, 2019). Among the many available DNN models, we selected this specific model because previous work has shown it to best approximate human metamers under indirect exploration paradigms (Feather et al.,, 2023). In our MAME framework, the model is used to define the direction of exploration. Therefore, in order to efficiently estimate human metamers under a limited number of conditions, it is preferable that the model be biologically plausible. In particular, we note that manipulating high-level features is more difficult in other models, as shown in the comparison results provided in the Supplementary Material (Fig. S1).
Next, we detail how the search space was concretely defined (Fig. 3). First, natural images were passed through the robust ResNet50, while extracting Gram matrices from the activations of each layer as position-invariant features. Here, we used Gram matrices because the present study explored metamers based on the equivalence of perceptual appearance and focused summary statistics like textures representations (Gatys et al.,, 2015, 2016; Jagadeesh and Gardner,, 2022).
Figure 3: Definition of exploration directions in the present study. An input image is passed through a robust ResNet50 model, and activations are extracted from the conv1, layer3, and avgpool layers. For each layer, filter-by-filter Gram matrices are computed and vectorized. Using these vectorized features from the ImageNet dataset, ICA is applied, and the top three components with the highest explained variance are selected as exploration axes for each layer.
A Gram matrix represents the similarity of activations across multiple filters, summed over spatial positions. Given the activation of the $l$-th layer of a CNN for an input image as $F_{jk}$, where $j$ denotes the filter index and $k$ denotes the spatial position, the Gram matrix $G_{ij}$ is defined as:
$G_{ij}=\sum_{k}F_{ik}F_{jk}.$ (1)
In this study, Gram matrices were extracted from multiple layers of the model using 1,000 images randomly selected from the ImageNet validation dataset (Deng et al.,, 2009).
After extracting the Gram matrix features, ICA was performed on them, and components with high explanatory power were selected, thereby considering images as points in the ICA coordinate space. The exploration direction was defined along the axes of ICA from the reference stimulus.
Although several dimensionality reduction methods could be considered, we used ICA for the following reasons. First, principal component analysis (PCA), which preserves orthogonality in the input space, is not suitable for disentangling high-level representations in DNNs (as described in Fig. S2). In addition, ICA standardizes components to have unit variance, making it easier to compare representations across different layers.
In the following, we describe in detail how dimensionality reduction of the metameric space was performed using ICA. ICA assumes non-Gaussianity of signals and computes a linear transformation $\mathbf{W}$ to separate observed signals $\mathbf{X}$ into independent components $\mathbf{S}$ as $\mathbf{S}=\mathbf{W}\mathbf{X}$. Each independent component (hereafter referred to as an ICA component) is standardized to have zero mean and a covariance matrix equal to the identity matrix. The Gram matrices of the target layers were vectorized to form feature vectors $\mathbf{x}$ for each image. Using the feature vector $\mathbf{x}$ of 1,000 images, an input data matrix $\mathbf{X}$ was constructed, and ICA was applied to obtain the independent component matrix $\mathbf{S}$:
$\mathbf{S}=\mathbf{W}\mathbf{X},\quad\mathbf{s}=\mathbf{W}\mathbf{x},\quad\mathbf{X}=\mathbf{S}\mathbf{A}^{\mathsf{T}}$ (2)
where $A$ represents the mixing matrix. The number of independent components was pre-set to 100, and the top 3 components with the highest contribution were selected. Practically, we used FastICA of the Python scikit-learn library to compute the components. The contribution of each component was quantified using the explained variance (EV) defined as:
$\mathrm{EV}_{i}=1-\frac{\|\mathbf{X}-\hat{\mathbf{X}}_{i}\|_{\text{F}}^{2}}{\|\mathbf{X}\|_{\text{F}}^{2}},$ (3)
where $\hat{X}_{i}$ is defined as:
$\hat{\mathbf{X}}_{i}=\mathbf{S_{i}}\mathbf{A}^{\mathsf{T}},$ (4)
with $\mathbf{S_{i}}$ being the matrix retaining only the $i$-th independent component. Unlike PCA, the total explained variance for multiple ICA components does not necessarily match the sum of the individual component contributions. The contribution of each component is shown in Fig. S4.
Finally, we computed the metameric space based on features extracted from the conv1, layer3, and avgpool layers of the robust ResNet50. These layers were selected to examine early, intermediate, and higher-level stages of visual processing, while avoiding layers affected by skip connections in the ResNet architecture. Both positive and negative directions of each ICA component were used in the exploration.
2.3 Generation of Target Images
This section describes the method for generating images at designated positions using gradient descent. The image modification process was performed by specifying the target layer (conv1, layer3, and avgpool), the ICA component of interest (components 0, 1, 2), the direction of change (positive or negative), the magnitude of change (a positive scalar corresponding to the target value), and the reference image (i.e., the initial image of the modification).
The image generation procedure is as follows:
  1. 1.
    Compute the ICA components of the reference image for the target layer as $\mathbf{y}_{\text{o}}=\mathbf{W}\mathbf{x}_{\text{o}}$, where $\mathbf{x}_{\text{o}}$ is the vectorized Gram matrix of the target layer.
  2. 2.
    Define the target component as $\mathrm{p}$ and the target value as $t$. The modified ICA component $\mathbf{y}_{\text{t}}$ is given by:
    $\mathbf{y}_{\text{t}}=\mathbf{y}_{\text{o}}+t\mathbf{e}_{\text{p}},$ (5)
    where $\mathbf{e}_{i}$ is a unit vector with 1 at the $i$-th component and 0 elsewhere. This operation modifies only the target component while keeping all others unchanged. If the change direction is negative, the unit vector is multiplied by 1-1.
  3. 3.
    In each iteration, compute the ICA components of the current image as $\mathbf{y}_{\text{c}}=\mathbf{W}\mathbf{x}_{\text{c}}$.
  4. 4.
    Compute the mean squared error (MSE) between $\mathbf{y}_{\text{t}}$ and $\mathbf{y}_{\text{c}}$ as the loss function.
  5. 5.
    Perform gradient descent using the Adam optimizer to minimize the loss and update pixel values of the current image while keeping the CNN model parameters fixed.
Examples of image generation for the three layers and three components used in the experiment are shown in Fig.4 and Fig. S3.
Figure 4: An example of the images used in the experiment, generated with three layers and three components each. The numbers (in red and blue) below the layer names represent the target ICA values used during image generation.The magnitude of modulation in this visualization is scaled according to the experimental results.
The selection of reference images for initialization was based on two criteria. First, we aimed to use images located near the zero point of the ICA components. Second, to prevent participants in the behavioral experiment from memorizing individual images, we needed to sample many images. For this purpose, from 50,000 images in the ImageNet validation dataset, we sampled images for three target layers whose ICA component distances from the origin were within the lowest 20%. Among these, 614 images that met the selection criteria across all three layers were used as experimental stimuli (Table 1).
Table 1: ICA component criteria and the number of selected images
Criterion* Layer Threshold Number of Img
Lowest 20% conv1 1.22 614
layer3 1.73
avgpool 0.79
* An image satisfies a given criterion if its Euclidean distance in each layer is below the corresponding threshold.
2.4 Feedback from Behavioral task
As feedback in the present experiment, we employed an ABX task to measure perceptual similarity. While various forms of feedback could be considered—such as neural similarity or categorical judgments—our goal was to demonstrate the effectiveness of the novel MAME framework under specific conditions. In general, measuring perceptual appearance psychophysically is more time-consuming than categorical judgment in one trial. In contrast, computing the equivalence of neural or BOLD activations in neurophysiological or fMRI studies is similar time course to experiments for perceptual appearance, although the experimental cost is much higher in neuroscience experiments. The ABX task was chosen because it offers a comparable trial duration to neural similarity experiments, while being less costly to implement. This makes it a suitable choice for assessing the feasibility of the MAME framework in an online experimental environment.
The experiment consisted of an ABX test comparing generated images with reference images. Based on the participants’ response data, each image was updated until the exploration of human metamer boundaries converged.
Eight participants took part in the experiment. The study was conducted with approval from the ethics committee of the research institution (detailed information will be disclosed after the double-blind peer-review process).
The MAME framework was implemented using a Python web framework (Django) and a browser-based experimental environment with JavaScript (Adolphe et al.,, 2022). Image generation was performed on a server machine with an NVIDIA RTX A2000 GPU, and the generated images were presented in the Google Chrome browser via a local server.
The experiment was conducted in a dark room, with participants’ heads fixed at a distance of 70 cm from the monitor. In all tasks, participants were instructed to fixate on a fixation point (Thaler et al.,, 2013) at the center of the monitor (Fig. 5).
Figure 5: Overview of the ABX test. A 200 ms stimulus presentation was alternated with a 500 ms blank period.
Furthermore, to ensure that participants were indeed fixating on the designated point during the experiment, gaze tracking was performed using the Tobii Pro Spark (Tobii, Sweden). Trials in which fixation was not maintained were not used for parameter updates.
The procedure of the ABX task is as follows (Fig. 5). In each trial, participants are presented with three stimuli in sequence: stimulus A, stimulus B, and stimulus X. The first two stimuli (A and B) are the reference and generated images in a random order. The third stimulus (X) is a test stimulus that is the same as A or B. Participants are required to identify whether X is the same as A or B. The presentation position was 4°, 8°, or 12° from the center depending on the eccentricity condition. The size of the presentation image was 4° x 4° in the visual angle.
Based on the results of each ABX test, the target value was updated using the staircase method (2-up 1-down rule). Specifically, if a participant correctly answered two consecutive trials in the ABX task, the boundary of the human metameric space was narrowed by one step. Conversely, if the participant made an error in a single trial, the boundary was shifted in the direction of expansion. Here, one step was defined as a fixed value for each layer, determined based on preliminary experiments: 10 for conv1, 0.3 for layer3, and 0.02 for avgpool. The threshold was estimated by averaging five reversals between correct and incorrect responses for each condition.
In summary, the experimental conditions consisted of three layers, three ICA components each explored in both positive and negative directions, and three levels of stimulus eccentricity (4°, 8°, and 12°). This resulted in a total of 54 conditions (3 × 3 × 2 × 3).
The experiment was conducted over two days for each participant. Each participant engaged in 15 experimental blocks, where each block consisted of 90 trials and took approximately 9–12 minutes to complete. In each block, the stimulus presentation position was fixed, and 18 conditions were presented five times each in a randomized order. Across the 15 blocks, three different stimulus presentation positions were used, with each position appearing in five blocks. The order of these positions was randomly determined at the start of the experiment. Following this protocol, each condition was tested 25 times using the ABX test.
3 Results
The MAME framework was applied to each subject, and their metameric boundaries were obtained as ICA component values. Furthermore, analyses based on the RMS contrast of images and Structural Similarity Index Measure (SSIM) were conducted, allowing the metameric boundaries to be computed using these indices.
3.1 Metameric Boundaries in ICA Components
The metameric boundaries represented by ICA components, obtained directly from the experiment, are shown in Table 2, Figure S5, and Figure S6. As the effects of the ICA component and direction conditions were small (Fig. S6), we first computed within-subject averages across these conditions. These mean values were averaged across different participants and shown in Table 2 and Figure S5.
Table 2: Mean threshold values computed by averaging across ICA components, search directions, and subjects for each layer and eccentricity. The standard deviation represents inter-subject variability, computed from the threshold values that were first averaged within each participant over components and directions. ($n=8$).
Layer Eccentricity (°) Mean Std
conv1 4 47.915 8.8166
8 63.0404 11.8613
12 63.7918 8.5024
layer3 4 0.486 0.1275
8 0.5983 0.0686
12 0.6407 0.0533
avgpool 4 0.1538 0.0182
8 0.153 0.0253
12 0.1688 0.0203
Across all subjects, the threshold values exhibited a logarithmic scale dependency on the layer. Specifically, the mean thresholds of conv1 were the highest in conv1, followed by layer3, and lowest in avgpool, decreasing as the layer depth increased. These results indicate that the boundaries of human metameric space are broader in the direction suggested by the robust CNN for lower-level features. This finding suggests a discrepancy in functional alignment between the model and human perception.
In contrast, the differences in thresholds due to variations in ICA components and direction were relatively minor compared to the inter-layer differences  2. This suggests that the human metameric space has uniform boundaries across different dimensional directions of the ICA space.
The results for eccentricity showed a gradual trend across conditions, with a particularly noticeable tendency for thresholds to increase as eccentricity increased in conv1 and layer3. To evaluate this trend in detail in terms of sensitivity to image contrast, the next section analyzes the results using RMS contrast and SSIM.
3.2 Metameric Boundaries Based on RMS Contrast and SSIM
The metameric boundaries obtained from the experiment were represented using the RMS contrast values of grayscale difference images. The RMS contrast is an index that quantifies the overall contrast strength of a luminance image and is defined as:
$C_{\mathrm{R}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(I_{i}-\langle\mathbf{I}\rangle)^{2}}$ (6)
where $N$ denotes the number of pixels, $I_{i}$ represents the luminance value of each pixel, and $\langle I\rangle$ is the mean luminance of the image.
In this analysis, for each layer and eccentricity, grayscale difference images were generated by subtracting the reference image from the perturbed image using the target ICA component values at the metameric boundary (Figure 6a). The RMS contrast of these difference images was then computed.
Figure 6: (a) Examples of threshold images for each layer (displayed in grayscale). The target value $t$ is determined based on the threshold for each layer. The rightmost column shows the difference images. (b) The RMS contrast of the difference image was computed. Images were generated with $n=30$ based on the average thresholds across participants from the behavioral experiment, and the mean and standard deviation of the RMS contrast of the difference image between the original and perturbed images are plotted. Lower values indicate smaller contrast differences. (c) Similarly, the SSIM between the original and perturbed images was computed, and the mean and standard deviation are plotted. Higher values indicate higher similarity between the original and perturbed images.
The RMS contrast values exhibited a linear relationship with the ICA target values. When visualizing the RMS contrast at the metameric boundary, it was observed that the values were particularly high in conv1 (Figure 6b). This suggests that, when measured using RMS contrast, human perception appears to be less sensitive to changes in conv1 compared to other layers.
Furthermore, the trend in eccentricity became more pronounced when analyzed using RMS contrast. Specifically, the slope of the threshold as a function of eccentricity was steepest in conv1, followed by layer3 and then avgpool, where it was more gradual. The finding that the effect of eccentricity diminishes at higher-level features has not been reported in previous studies and represents a novel contribution from the perspective of cognitive neuroscience.
Furthermore, these trends are not limited to comparisons based on RMS contrast, which reflects a global image statistic. We also evaluated perceptual sensitivity using the Structural Similarity Index (SSIM), which takes into account structural similarity between two images (Figure 6c). SSIM assesses local structural similarity within a sliding window. Consistent with the results based on RMS contrast, metameric images generated from the conv1 layer showed lower sensitivity, and the influence of eccentricity became progressively weaker as higher-level features were used.
4 Discussion
4.1 Summary of Main Findings
The present study proposed the MAME framework, an experimental framework to directly explore human metamers using online feedback by humans with weak supervision of artificial models. The effectiveness of this framework was demonstrated by successfully estimating multi-dimensional image metamer boundaries. Experimental results indicated that, among image generation layers, the threshold in conv1 was higher based on the magnitude of ICA thresholds, RMS contrast, and SSIM, suggesting that changes in this layer were less noticeable compared to others. Furthermore, the effect of eccentricity on sensitivity was less pronounced for higher-level features.
To contextualize our findings, it is important to clarify what model metamers represent in the current study. In our MAME framework, when an image is shifted even slightly in the model’s exploration direction from the reference image, the shifted image is no longer a model metamer for the model. This is because exploration is carried out in the model’s activation space, where model metamers are defined as stimuli that are distinct in image space but equivalent in the target activation space. As a result, a single step in this space exceeds the model’s metameric boundary in image space.
This relationship implies that, in the current exploration paradigm, lower mean thresholds observed in human responses at a given hierarchical level indicate a closer alignment between the metamer boundaries of the human and the model. In this regard, the low threshold found in the avgpool layer suggests a closer functional alignment between the metameric spaces of humans and the model.
4.2 Alignments of Higher and Lower Visual Processing in Humans and Models
The finding related to the alignment of higher image features is consistent with recent studies using robust CNNs (Feather et al.,, 2019, 2023; Gaziv et al.,, 2024). For instance, it has been shown that adversarial attacks on high-level activations can induce category modulations that are also perceptible to humans (Gaziv et al.,, 2024). Furthermore, the alignment of high-level features has been a topic of intense discussion in both visual and LLM literature. For example, the Platonic Representation Hypothesis (Huh et al.,, 2024) proposes that ’neural networks, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces.’ From this perspective, the strong alignment of high-level representations between humans and modern neural models, trained with various data, might be a straightforward consequence.
However, our experimental results also suggest a functional misalignment between models and humans in low-level processing. Given that convolutional neural networks (CNNs) are originally inspired by the information processing of simple and complex cells in the primary visual cortex (Fukushima,, 1980), this result may seem counterintuitive. One reason is that the first-layer filters in ResNet50 have a high-resolution size of 7×7 pixels and lack low-frequency components. Additionally, while the correlation between filters is known to be important for visual processing (Freeman and Simoncelli,, 2011; Okazawa et al.,, 2015), not all correlations among early-layer features in CNNs (i.e., the Gram matrix) might be necessarily relevant. Some attempts to address such early-stage mismatches have been reported to contribute to the development of more generalizable models (Dapello et al.,, 2020). Considering our finding that higher-level features are less affected by eccentricity, reconsidering the early-stage architecture of models might be crucial for improving functional alignment between humans and models in the future.
4.3 Limitations and Broader Future Applications
The main limitation of our study is that the explored stimulus conditions in our experiment do not comprehensively cover all possible directions of human metamer space in terms of clarifying human functional processing. While we used Gram matrices to define metameric equivalence, activation differences could also be used. Conducting complementary experiments with such alternative conditions, involving a larger set of models and participants, will be essential for further advancing our understanding of brain function.
Finally, it is noteworthy to discuss the generalizability of the proposed MAME framework. Since the framework can be executed in a web browser, it is free from hardware constraints. It can be applied even in environments with strong magnetic fields, such as fMRI, and is also suitable for crowdsourcing experiments. These advantages position MAME as a foundational tool for a wide range of applied research.
5 Conclusions
In this study, we proposed the Multidimensional Adaptive Metamer Exploration (MAME) framework, which combines hierarchical neural network–based image generation with human-in-the-loop exploration, to directly map the multidimensional human metameric space. Our evaluation experiment demonstrated that MAME can measure this space within a single psychophysical experiment. The results further revealed that a recent human-aligned model exhibits a substantial functional mismatch with human processing for low-level features, despite closer alignment at higher representational stages. Because the user interface of MAME operates entirely within a web browser, it can be applied in fMRI environments and large-scale crowdsourcing experiments, making it a diverse tool for exploring metamers in human vision and for advancing interpretable AI.
6 Acknowledgments
We used ChatGPT (OpenAI, GPT-4) for English language editing. All content was reviewed and verified by the authors.